# A Measurement of the Largest-Scale CMB $E$-mode Polarization with CLASS

### v1

@@ 5.1 Simulations and Noise Covariance Matrices
Low-resolution noise covariance matrices are needed for optimal foreground reduction and power spectrum estimation. For Planck PR4 30 GHz maps and WMAP maps, we used their official data product; for the SRoll2 version of the Planck maps, we followed the approach in P20 and used the suboptimal choice of noise covariance matrices built for Planck PR3 products (FFP8; Planck Collaboration XII, 2016). CLASS noise covariance matrix was built from noise-only simulations that ignore the systematics introduced by map non-linearity. All noise covariance matrices were augmented by a $20\,\mathrm{nK}$ regularization white noise at $N_{\mathrm{side}}=16$. As with the maps, when the external noise covariance matrices were used with the CLASS map, they were filtered with the transfer matrix $\mathbf{F}_{16,16}$.
@@ 5.2 Foreground Reduction
In addition to the CLASS 90 GHz, we chose the SRoll2 processing of the Planck HFI 100/143 GHz maps (Delouis et al., 2019) as the CMB channel ($\mathbf{m}_{c}$) for cross-correlation. Both Planck and CLASS maps need to be cleaned of synchrotron and dust polarization emission. Two data splits (ds) of the SRoll2 353 GHz map were used as dust templates ($\mathbf{m}_{d}$) for CLASS and Planck maps. Due to the lack of a consistent LFI processing from the SRoll2 data release (Delouis et al., 2019), we used the WMAP $K$-band (Bennett et al., 2013) and the Planck PR4 version of the 30 GHz map (Planck Collaboration Int. LVII, 2020) as synchrotron templates ($\mathbf{m}_{s}$) for CLASS and Planck maps respectively. We did not use the previously released CLASS 40 GHz map (L23; Eimer et al. 2024) as the synchrotron tracer mainly because of its lower sensitivity to synchrotron signal compared to lower frequency tracers. A joint analysis of CLASS multi-frequency maps is planned for future work.
@@ 5.3 $-\ell$ CMB Spectrum
Figure 12 shows the per-multipole uncertainty from the noise (i.e., not including sample variance, black) and the total error budget (blue, including the sample variance from the fiducial model). Figure 13 shows the covariance matrix of the estimate spectra at $2\leq\ell\leq 10$ among all three spectra: $EE$, $BB$, and $EB$. The smallness of the off-diagonal correlation coefficients suggests that the transfer matrix treatment has largely corrected for mode coupling and polarization leakage due to filtering and the cut sky.
Figure 13: Transfer matrix-corrected xQML spectra correlation matrix ($2\leq\ell\leq 10$) for CLASS $\times$ Planck over the f050 mask. The matrix is built from 500 simulations combining the foreground-cleaned simulation suite with signals drawn from the fiducial $\Lambda$CDM model. The correlation amplitudes are expressed as percentages and labeled in the lower triangle of the matrix. Red and blue colors indicate positive and negative correlations, respectively. The diagonal elements are colored black.
As an additional check, we tested the pipeline for a CLASS null map by differencing two map splits and cross-correlating with the foreground-reduced Planck coaddition map. The cross-spectra, denoted as “Null $\times$ Planck”, were compared to an ensemble of simulations. The simulations for the Planck map contained CMB realizations ($\tau=0.054$) and foreground cleaning residuals as above, but only instrumental noise was included for the CLASS null map. The per-multipole PTE values are tabulated in Table 5.3 (first three columns). The most extreme values are found to be $0.002/0.998$ in the $EB$ spectrum, approaching the precision limited by the number of simulations. Considering the total number of tests conducted, these results indicate good overall consistency.
Table 3: PTE values for the Null $\times$ Planck and CLASS $\times$ Planck cross spectra for each multipole $\ell$ and every polarization spectrum. In the former case, the null spectra are compared to the null model, and in the latter case, the comparison is made against the best-fit model with $\tau=0.053$ (Section 5.3). Given the limited number of simulations, PTE values are calculated from the cumulative distribution of the simulations using kernel density estimation. The last row lists the one-tailed p-value from a KS test assessing the uniformity of PTE values across multipoles. The color scheme follows that used in Table 4.1.
@@ 5.4 Simulation-based $\tau$ Inference
in good agreement with Planck-only results (P20; Tristram et al. 2024). For the best-fit cosmology, we computed the PTE values for each multipole and spectrum by comparing the data estimate to the distribution of the 500 signal-plus-noise simulations, and the results are listed in Table 5.3 (last three columns). Overall, the PTE values are consistent with uniform distributions between 00 and $1$ across multipoles, as confirmed by the one-tailed KS test for $EE$/$BB$ and $EB$ spectra, suggesting that the model is a good fit to the data. Since there is no reionization signal present in the $EB$ and $BB$ spectra (assuming no primordial tensor mode) and the CMB lensing signal is negligible, the check on the $B$-mode spectra serves as an additional null test for the analysis pipeline. Building the covariance matrix for the $EE$ power spectrum over $2\leq\ell\leq 30$ from 500 simulations at the best-fit cosmology with appropriate bias correction (Hartlap et al., 2007) yielded $\chi^{2}=32.7$ for $28$ degrees of freedom, corresponding to a PTE of $0.22$. Repeating the likelihood analysis but replacing the CLASS data with simulations fixing $\tau=0$, we found that the best-fit result from CLASS rejects the null hypothesis of no reionization at $99.4\%$ significance.
@@ 5.5 Forecast
Since the retained power at the lowest $\ell$ in the current analysis is essentially zero, $\lambda^{2}$ is approximately the improved transfer function at $\ell=2$. The projected map depth improvement is achieved by scaling the noise-only simulations and noise covariance matrices by constant factors. This implies an optimistic assumption that the $-\ell$ red noise scales well with the map depth and does not inflate as the time-domain filtering is relaxed (before applying the transfer function correction). These assumptions might not hold for components such as linearly-polarized atmospheric emission (Li et al., 2023a; Coerver et al., 2024).

### v2

@@ 5.1 Simulations and Noise Covariance Matrices
Low-resolution noise covariance matrices are needed for optimal foreground reduction and power spectrum estimation. For Planck PR4 30 GHz maps and WMAP maps, we used their official data product; for the SRoll2 version of the Planck maps, we followed the approach in P20 and used the suboptimal choice of noise covariance matrices built for Planck PR3 products (FFP8; Planck Collaboration XII, 2016). CLASS noise covariance matrix was built from noise-only simulations that ignore the systematics introduced by map non-linearity. All noise covariance matrices were augmented by a $20\,\mathrm{nK}$ regularization white noise at $N_{\mathrm{side}}=16$ to ensure numerical stability during matrix inversion. As with the maps, when the external noise covariance matrices were used with the CLASS map, they were filtered with the transfer matrix $\mathbf{F}_{16,16}$.
@@ 5.2 Foreground Reduction
In addition to the CLASS 90 GHz, we chose the SRoll2 processing of the Planck HFI 100/143 GHz maps (Delouis et al., 2019) as the CMB channel ($\mathbf{m}_{c}$) for cross-correlation. Both Planck and CLASS maps need to be cleaned of synchrotron and dust polarization emission. Two data splits (ds) of the SRoll2 353 GHz map were used as dust templates ($\mathbf{m}_{d}$) for CLASS and Planck maps. Due to the lack of a consistent LFI processing from the SRoll2 data release (Delouis et al., 2019), we used the WMAP $K$-band (Bennett et al., 2013) and the Planck PR4 version of the 30 GHz map (Planck Collaboration Int. LVII, 2020) as synchrotron templates ($\mathbf{m}_{s}$) for CLASS and Planck maps respectively. We did not use the previously released CLASS 40 GHz map (L23; Eimer et al. 2024) as the synchrotron tracer mainly because of its lower sensitivity to synchrotron signal compared to lower frequency tracers and the complexities involved in accounting for different transfer functions in CLASS maps. A map-level joint analysis of CLASS multi-frequency maps with WMAP/Planck to maximize sensitivity is planned for future work.
@@ 5.3 $-\ell$ CMB Spectrum
Figure 13 shows the per-multipole uncertainty from the noise (i.e., not including sample variance, black) and the total error budget (blue, including the sample variance from the fiducial model). Figure 14 shows the correlation matrix of the estimate spectra at $2\leq\ell\leq 10$ among all three spectra: $EE$, $BB$, and $EB$. Given the ensemble size of 500, the off-diagonal elements are expected to have statistical fluctuations at $5\%$ level (Lueker et al., 2010). The dominant correlations appear on the second diagonal, consistent with the performance of xQML over the cut sky (Vanneste et al., 2018). Other significant correlation coefficients are associated with the data and depend on data combinations. The overall smallness of the off-diagonal correlations suggests that the transfer matrix treatment has largely corrected for mode coupling and polarization leakage due to filtering and the cut sky.
Figure 14: Transfer matrix-corrected xQML spectra correlation matrix ($2\leq\ell\leq 10$) for CLASS $\times$ Planck over the f050 mask. The matrix is built from 500 simulations combining the foreground-cleaned simulation suite with signals drawn from the fiducial $\Lambda$CDM model. The correlation amplitudes are expressed as percentages and labeled in the lower triangle of the matrix. Red and blue colors indicate positive and negative correlations, respectively. The diagonal elements are colored black. The correlation coefficient amplitudes peak around 20% on the second off-diagonals, primarily due to the inherent mode coupling from the xQML estimator (Vanneste et al., 2018).
As an additional check, we tested the pipeline for a CLASS null map by differencing two map splits and cross-correlating with the foreground-reduced Planck coaddition map. The cross-spectra, denoted as “Null $\times$ Planck”, were compared to an ensemble of simulations. The simulations for the Planck map contained CMB realizations ($\tau=0.054$) and foreground cleaning residuals as above, but only instrumental noise was included for the CLASS null map. The per-multipole PTE values are tabulated in Figure 15 (first three columns). The most extreme values are found to be $0.002/0.998$ in the $EB$ spectrum, approaching the precision limited by the number of simulations. Considering the total number of tests conducted, these results indicate good overall consistency.
Figure 15: PTE values for the Null $\times$ Planck and CLASS $\times$ Planck cross spectra for each multipole $\ell$ and every polarization spectrum. In the former case, the null spectra are compared to the null model, and in the latter case, the comparison is made against the best-fit model with $\tau=0.053$ (Section 5.3). Given the limited number of simulations, PTE values are calculated from the cumulative distribution of the simulations using kernel density estimation. The last row lists the one-tailed p-value from a KS test assessing the uniformity of PTE values across multipoles. The color scheme follows that used in Figure 9.
@@ 5.4 Simulation-based $\tau$ Inference
in good agreement with Planck-only results (P20; Tristram et al. 2024). For the best-fit cosmology, we computed the PTE values for each multipole and spectrum by comparing the data estimate to the distribution of the 500 signal-plus-noise simulations, and the results are listed in Figure 15 (last three columns). Overall, the PTE values are consistent with uniform distributions between 00 and $1$ across multipoles, as confirmed by the one-tailed KS test for $EE$/$BB$ and $EB$ spectra, suggesting that the model is a good fit to the data. Since there is no reionization signal present in the $EB$ and $BB$ spectra (assuming no primordial tensor mode) and the CMB lensing signal is negligible, the check on the $B$-mode spectra serves as an additional null test for the analysis pipeline. Building the covariance matrix for the $EE$ power spectrum over $2\leq\ell\leq 30$ from 500 simulations at the best-fit cosmology with appropriate bias correction (Hartlap et al., 2007) yielded $\chi^{2}=32.7$ for $28$ degrees of freedom, corresponding to a PTE of $0.22$. Repeating the likelihood analysis but replacing the CLASS data with simulations fixing $\tau=0$, we found that the best-fit result from CLASS rejects the null hypothesis of no reionization at $99.4\%$ significance.
@@ 5.5 Forecast
Since the retained power at the lowest $\ell$ in the current analysis is essentially zero, $\lambda^{2}$ is approximately the improved transfer function at $\ell=2$. The projected map depth improvement is achieved by scaling the noise-only simulations and noise covariance matrices by constant factors. This implies an optimistic assumption that the $-\ell$ red noise scales well with the map depth and does not inflate as the time-domain filtering is relaxed (before applying the transfer function correction). These assumptions might not hold for components such as linearly-polarized atmospheric emission (Li et al., 2023a; Coerver et al., 2025).